# Query-Efficient Zeroth-Order Algorithms for Nonconvex Optimization

### v1

@@ Abstract
Zeroth-order optimization (ZO) has been a powerful framework for solving black-box problems, which estimates gradients using zeroth-order data to update variables iteratively. The practical applicability of ZO critically depends on the efficiency of single-step gradient estimation and the overall query complexity. However, existing ZO algorithms cannot achieve efficiency on both simultaneously. In this work, we consider a general constrained optimization model with black-box objective and constraint functions. To solve it, we propose novel algorithms that can achieve the state-of-the-art overall query complexity bound of $\mathcal{O}(d/\epsilon^{4})$ to find an $\epsilon$-stationary solution ($d$ is the dimension of variable space), while reducing the queries for estimating a single-step gradient from $\mathcal{O}(d)$ to $\mathcal{O}(1)$. Specifically, we integrate block updates with gradient descent ascent and a block gradient estimator, which leads to two algorithms, ZOB-GDA and ZOB-SGDA, respectively. Instead of constructing full gradients, they estimate only partial gradients along random blocks of dimensions, where the adjustable block sizes enable high single-step efficiency without sacrificing convergence guarantees. Our theoretical results establish the finite-sample convergence of the proposed algorithms for nonconvex optimization. Finally, numerical experiments on a practical problem demonstrate that our algorithms require over ten times fewer queries than existing methods.
@@ 1  Introduction
where $h:\mathbb{R}^{d_{x}}\to\mathbb{R}$ is the objective function and each $c_{j}:\mathbb{R}^{d_{x}}\to\mathbb{R},\forall j\in\mathcal{J}$ is a constraint function. Both $h(x)$ and $c_{j}(x)$ do not have analytical expressions and are treated as black boxes, i.e., only the input $x$ and the corresponding deterministic function outputs $h(x)$ or $c_{j}(x)$ are observable. Neither $h$ nor $c_{j},\forall j\in\mathcal{J}$ is necessarily convex.
@@
Problems in the form of (1) arise across many domains, such as power systems (Hu et al.,, 2024; Zhou et al.,, 2025), simulation optimization (Park and Kim,, 2015), and machine learning (Nguyen and Balasubramanian,, 2023). However, traditional model-based or gradient-based algorithms are inapplicable to problem (1), as they rely on first-order or second-order information (e.g., gradients or Hessians) of $h(x)$ and $c_{j}(x)$, which is not available. Zeroth-order optimization (ZO), a representative method in derivative-free optimization, offers a promising approach to this type of optimization problem, and has been broadly applied (Fu et al.,, 2015; Liu et al., 2020a, ; Malladi et al.,, 2023; Lam and Zhang,, 2024). The fundamental idea behind ZO is to construct estimators of first-order information using zeroth-order data (Berahas et al.,, 2022), and integrate these estimators into gradient-based algorithms, such as gradient descent, to seek optimal or high-quality solutions.
@@
Under the iterative ZO framework, the efficiency of single-step gradient estimation and overall query complexity jointly determine the practical applicability of ZO (Scheinberg,, 2022). They refer to the number of function values required to generate a single-step gradient and a final solution, respectively. The traditional coordinate-wise gradient estimation (CGE) requires estimating partial gradients along all dimensions separately based on finite differences of function values (Kiefer and Wolfowitz,, 1952). Although CGE-based algorithms can generally enjoy state-of-the-art overall query complexities due to the controllable bias and variance of CGEs (Xu et al.,, 2024; Zhou et al.,, 2025), the requirement of $\mathcal{O}(d)$ queries for estimating a single-step gradient makes them inefficient for high-dimensional problems ($d$ is the dimension of variable space). In contrast, the prevalent randomized gradient estimation (RGE) only requires one or two function values to construct a gradient estimator along a random direction (Flaxman et al.,, 2005; Nesterov and Spokoiny,, 2017). RGE-based algorithms have demonstrated excellent performance in unconstrained problems. However, they suffer from slow convergence when applied to constrained cases (such as (1)) due to large variances of gradient estimation (Liu et al., 2020a, ). This significantly limits their practical performance. In view of this dilemma, a fundamental question arises:
To solve (1), can we design zeroth-order algorithms that are query-efficient regarding both single-step and overall complexities?
@@
In this paper, we provide a positive answer to this question. We will utilize the framework of random block updates to design novel and query-efficient ZO algorithms for solving problem (1), and show that the proposed algorithms enjoy controllable single-step efficiency and the best-known overall query complexities.
@@ 1.1  Main Contributions
We assume simultaneous zeroth-order access to $h(x)$ and $c_{j}(x),\forall j\in\mathcal{J}$ (i.e., we can observe all the function evaluations of $h(x)$ and $c_{j}(x),\forall j\in\mathcal{J}$ simultaneously via querying a $x$) but no gradient information. To handle the black-box constraints in (1), we adopt a primal-dual framework by reformulating it as a deterministic min-max problem:
@@
where $f(x,y)=h(x)+y^{\text{T}}c(x)$ is the Lagrange function of problem (1). Wherein, $c(x)=(c_{1}(x),\cdots,c_{d_{y}}(x))^{\text{T}}$ with $d_{y}=|\mathcal{J}|$; $\mathcal{Y}=\{y\in\mathbb{R}^{d_{y}}|y\geq 0\}$ is the feasible set of Lagrange multiplier. Clearly, $f(x,y)$ is nonconvex-concave, i.e., nonconvex in $x$ and concave in $y$, when $h(x)$ and $c_{j}(x),\forall j\in\mathcal{J}$ are not assumed convex. Then, solving problem (2) can provide optimal or high-quality solutions to problem (1) (Nesterov et al.,, 2018). The detailed contributions of this work are summarized as follows.
Table 1: Comparison of single-step and overall query complexities
@@
SZO-ConEX (Nguyen and Balasubramanian,, 2023) RGE $\mathcal{O}(1)$ $\mathcal{O}(d/\epsilon^{6})$
ZOAGP (Xu et al.,, 2024) CGE $\mathcal{O}(d)$ $\mathcal{O}(d/\epsilon^{4})$
@@
Note: In our algorithms, $d=d_{x}$. $b$ is the block size that can be chosen from $\{1,2,\cdots,d\}$. Overall queries refer to the number of queries required to achieve an $\epsilon$-stationary/KKT point.
@@
In this research, we adopt the widely-used gradient descent ascent (GDA) framework to solve problem (1). However, directly applying RGE or CGE in GDA cannot exhibit satisfactory performance in both single-step and overall complexity (see our detailed discussion in Section 2.2). To address this, we combine the framework of block updates with GDA and smoothed GDA to develop two novel algorithms, called zeroth-order block gradient descent ascent (ZOB-GDA) and zeroth-order block smoothed gradient descent ascent (ZOB-SGDA). Rather than estimating a full gradient at each step, they randomly select a block of coordinates and update the variables using block coordinate-wise gradient estimations (BCGEs). The adoption of the BCGEs effectively controls the bias and variance of gradient estimations to be negligible and thereby accelerates convergence. Moreover, the block size is adjustable to control the number of queries required to construct a single-step gradient.
Best-Known Query Complexities with Controllable Single-Step Efficiency.
We establish finite-sample guarantees for the proposed algorithms by analyzing the min-max problems (2) in nonconvex-concave settings. The query complexity results are summarized in Table 1 and compared with two representative algorithms. Specifically, ZOB-GDA can find an $\epsilon$-stationary point of $f(x,y)$ with a query complexity bound $\mathcal{O}(d/\epsilon^{6})$, which differs from the bound for first-order GDA only by a factor of $d$. Moreover, ZOB-SGDA is shown to have the query complexity bound $\mathcal{O}(d/\epsilon^{4})$, which aligns with the best-known results for solving deterministic nonconvex-concave problems. Different from existing methods, our algorithms also benefit from controllable efficiency in single-step gradient estimation, which makes them query-efficient for both single-step and overall complexities. The numerical results demonstrate that our algorithms can require over 10 times fewer queries for both a single step and overall complexity compared to existing methods.
@@ 1.2  Related Work
Here, we provide a detailed discussion on prior related work in ZO, zeroth-order gradient descent ascent, and coordinate/block updates in ZO.
@@
ZO has emerged as a prevalent tool to solve black-box problems and found application across machine learning (Liu et al., 2020a, ; Nguyen and Balasubramanian,, 2023), power systems (Hu et al.,, 2024; Zhou et al.,, 2025), simulation optimization (Fu et al.,, 2015; Lam and Zhang,, 2024), large language models (Malladi et al.,, 2023; Zhang et al.,, 2024), etc. ZO originates from the stochastic approximation method in Kiefer and Wolfowitz, (1952), where CGE is applied to estimate partial gradients along all dimensions via finite differences of function values. This is inefficient for high-dimensional problems, even if parallel techniques can be applied (Scheinberg,, 2022). To improve single-step efficiency, one-point and two-point RGEs have been developed by estimating gradients along randomized directions (Flaxman et al.,, 2005; Nesterov and Stich,, 2017; Lam and Zhang,, 2024). Generally, RGE-based algorithms can achieve the same oracle complexities as their first-order counterparts in unconstrained problems, differing by a dimension-dependent factor (Liu et al., 2020a, ). However, RGEs suffer from large variance in gradient estimation in constrained problems (see Section 2.2 for a detailed discussion). Moreover, most literature considers simple constraints on the input $x$ that can be dealt with by projection operations (Duchi et al.,, 2015; Yuan et al.,, 2015; Jin et al.,, 2023; He et al.,, 2024), which cannot be used to solve problem (1).
@@
GDA is a classical framework for solving min-max problems and has been extensively studied (Nemirovski,, 2004; Nedić and Ozdaglar,, 2009; Lin et al.,, 2020; Xu et al.,, 2023; Zhang et al.,, 2020). It is also well-established and widely applied to solve zeroth-order min-max problems of the form (2) (Hu et al.,, 2024; Nguyen and Balasubramanian,, 2023). The authors of Liu et al., (2018) applied the two-point RGE to solve a composite optimization problem. Then, the standard zeroth-order GDA was applied to the general min-max problems (Liu et al., 2020b, ; Wang et al.,, 2023), while only the query complexity of nonconvex-strongly concave cases was established. Several variants of zeroth-order GDA have been developed for convex-concave settings, such as zeroth-order OGDA-RR (Maheshwari et al.,, 2022) and zeroth-order extra-gradient (Zhou et al.,, 2025), which can achieve the query complexity bounds of $\mathcal{O}(d^{4}/\epsilon^{2})$ and $\mathcal{O}(d/\epsilon^{2})$ to reach an $\epsilon$-optimal solution, respectively. For nonconvex–concave problems, Xu et al., (2024) proposed combining alternating gradient projection with CGEs to solve a min-max problem with the query complexity bound of $\mathcal{O}(d/\epsilon^{4})$ to obtain an $\epsilon$-stationary point. In Nguyen and Balasubramanian, (2023), the SZO-ConEX algorithm was designed based on RGEs to achieve the query complexity of $\mathcal{O}(d/\epsilon^{6})$ to derive an $\epsilon$-critical Karush-Kuhn-Tucker (KKT) point of the problem (1).
@@
The framework of coordinate/block updates is widely adopted in first-order optimization (Nesterov and Stich,, 2017; Latafat et al.,, 2019). The core idea is to apply the partial gradients along a subset of full dimensions to update the iterates. The applications of coordinate/block updates in ZO mainly lie in unconstrained problems (Lian et al.,, 2016; Cai et al.,, 2021), where only the coordinate/block gradients along a subset are estimated using coordinate/block CGEs or RGEs at each step. Their extension to constrained problems, however, remains relatively underexplored. In Shanbhag and Yousefian, (2021), the RGE was combined with zeroth-order block updates and projected gradient descent to solve a stochastic constrained problem. Moreover, in He et al., (2024) and Jin et al., (2023), a cyclic zeroth-order block coordinate descent method and a randomized zeroth-order coordinate descent method were proposed, respectively, to solve the deterministic constrained problems and achieve complexity bounds proportional to $\epsilon^{-2}$ for nonconvex optimization. However, all these methods require the constraint set to be projection-friendly and coordinate/block-structured, which is usually too restrictive in practical problems and inapplicable to non-analytical constraint sets (such as in our problem (1)).
@@ 2  Preliminaries
Notations.
For a positive integer $n$, we denote $n]:=\{1,2,\cdots,n\}$. For a vector $x\in\mathbb{R}^{d_{x}}$, denote $x(i)$ as its $i$th entry. For a differentiable function $h(x):\mathbb{R}^{d_{x}}\to\mathbb{R}$, denote $\nabla h(x)$ as its gradient at $x$ and $\nabla_{i}h(x),i\in[d_{x}]$ as the partial gradient along the $i$th dimension. Similarly, for a differentiable function $f(x,y):\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$, denote the partial gradient w.r.t. $x$ (and $y$) by $\nabla_{x}f(x,y)$ (and $\nabla_{y}f(x,y)$). Without further specification, $\|\cdot\|$ denotes the $\ell_{2}$-norm in Euclidean space. The Euclidean projection operator onto a closed convex set $\mathcal{X}$ is denoted by $\mathcal{P}_{\mathcal{X}}[\cdot]$.

### v2

# Query-Efficient Zeroth-Order Algorithms for Nonconvex Constrained Optimization

@@ Abstract
Zeroth-order optimization (ZO) has been a powerful framework for solving black-box problems, which estimates gradients using zeroth-order data to update variables iteratively. The practical applicability of ZO critically depends on the efficiency of single-step gradient estimation and the overall query complexities. However, existing constrained ZO algorithms cannot achieve efficiency on both simultaneously. In this work, we consider a general constrained optimization model with black-box objective and constraint functions. To solve it, we propose novel algorithms that can achieve the best-known overall query complexity bound of $\mathcal{O}(d/\epsilon^{4})$ to find an $\epsilon$-stationary solution ($d$ is the dimension of variables), while reducing the queries for estimating a single-step gradient from $\mathcal{O}(d)$ to $\mathcal{O}(1)$. Specifically, we integrate block gradient estimators with gradient descent ascent, which leads to two algorithms, ZOB-GDA and ZOB-SGDA, respectively. Instead of constructing full gradients, they estimate only partial gradients along random blocks of dimensions, where the adjustable block sizes enable high single-step efficiency without sacrificing convergence guarantees. Our theoretical results establish the finite-sample convergence of the proposed algorithms for nonconvex optimization. Finally, numerical experiments demonstrate the superior performance of our algorithms compared to existing methods.
@@ 1  Introduction
where $h:\mathbb{R}^{d_{x}}\to\mathbb{R}$ is the objective function and each $c_{j}:\mathbb{R}^{d_{x}}\to\mathbb{R},\forall j\in\mathcal{J}$ is a constraint function. In our problem, we have the variable dimension $d=d_{x}$. Both $h(x)$ and $c_{j}(x)$ do not have analytical expressions and are treated as black boxes, i.e., only the input $x$ and the corresponding deterministic function outputs $h(x)$ or $c_{j}(x)$ are observable. Neither $h$ nor $c_{j},\forall j\in\mathcal{J}$ is necessarily convex.
@@
Problems of the form (1) arise in many domains, such as power systems (Hu et al., 2024; Zhou et al., 2025), simulation optimization (Park and Kim, 2015), and machine learning (Nguyen and Balasubramanian, 2023). However, traditional model-based or gradient-based algorithms are generally not applicable to problem (1), as they rely on first-order or second-order information (e.g., gradients or Hessians) of $h(x)$ and $c_{j}(x)$ that is not available. Zeroth-order optimization (ZO), a representative method in derivative-free optimization, offers a promising approach to this type of optimization problem, and has been broadly applied (Fu and others, 2015; Liu et al., 2020a; Malladi et al., 2023; Lam and Zhang, 2025). The fundamental idea behind ZO is to construct estimators of first-order information using zeroth-order data (Berahas et al., 2022), and integrate these estimators into gradient-based algorithms, such as gradient descent, to seek optimal or high-quality solutions.
@@
Under the iterative ZO framework, the efficiency of single-step gradient estimation and overall query complexity jointly determine the practical applicability of ZO. These two metrics refer to the number of function values required to generate a gradient estimator and a final solution, respectively. The traditional coordinate-wise gradient estimation (CGE) requires estimating partial gradients along all dimensions separately based on finite differences of function values (Kiefer and Wolfowitz, 1952). CGE-based algorithms can generally enjoy state-of-the-art overall query complexity bounds due to the controllable bias and variance of CGEs (Xu et al., 2024). However, each CGE costs $\mathcal{O}(d)$ function evaluations ($d$ is the dimension of the variable space), which becomes inefficient in high-dimensional problems, even if parallel computing techniques are available (Scheinberg, 2022; Hu and Fu, 2025). In contrast, the prevalent randomized gradient estimation (RGE) only requires one or two function values to construct a gradient estimator along a random direction (Flaxman et al., 2005; Nesterov and Spokoiny, 2017). RGE-based algorithms have demonstrated excellent performance in unconstrained problems. However, they suffer from slow convergence when applied to constrained cases (such as (1)) due to the large variance of gradient estimations (Liu et al., 2020a). Existing methods must trade off low single-step efficiency against good overall complexity, and none of these methods is query-efficient in both for high-dimensional constrained problems.
@@
In view of this challenge, a fundamental question arises: To solve (1), is it possible to achieve the best-known query complexity bound while improving the single-step efficiency? Similar challenges have also been recognized in Scheinberg (2022) and Hu and Fu (2025). In this study, we provide a positive answer to this question by utilizing the framework of random block updates to design query-efficient ZO algorithms for solving (1). We show that our method can enjoy improved single-step efficiency and the best-known overall query complexity bound.
@@ 1.1  Main Contributions
We assume an oracle that returns $h(x)$ and $c_{j}(x),\forall j\in\mathcal{J}$ (i.e., we can observe all the function evaluations of $h(x)$ and $c_{j}(x),\forall j\in\mathcal{J}$ simultaneously via querying a $x$) but no gradient information. To handle the black-box constraints in (1), we adopt a primal-dual framework by reformulating it as a deterministic min-max problem:
@@
where $f(x,y)=h(x)+y^{\text{T}}c(x)$ is the Lagrange function of problem (1). Here, $c(x)=(c_{1}(x),\cdots,c_{d_{y}}(x))^{\text{T}}$ with $d_{y}=|\mathcal{J}|$; $\mathcal{Y}$ is the feasible set of Lagrange multipliers. Clearly, $f(x,y)$ is nonconvex-concave, i.e., nonconvex in $x$ and concave in $y$, when $h(x)$ and $c_{j}(x),\forall j\in\mathcal{J}$ are not assumed convex. Then, solving problem (2) can provide high-quality solutions to problem (1) (Nesterov, 2018). The detailed contributions of this work are summarized as follows.
Table 1: Comparison of single-step and overall query complexity bounds
@@
SZO-ConEx (Nguyen and Balasubramanian, 2023) RGE $\mathcal{O}(1)$ $\mathcal{O}(d/\epsilon^{6})$
ZOAGP (Xu et al., 2024) CGE $\mathcal{O}(d)$ $\mathcal{O}(d/\epsilon^{4})$
@@
Note: In our algorithms, $d=d_{x}$. $b$ is the block size that can be chosen from $\{1,2,\ldots,d\}$. Overall queries refer to the theoretical upper bound on the number of oracle queries required to achieve an $\epsilon$-stationary point.
@@
In this research, we adopt the widely-used gradient descent ascent (GDA) framework to solve problem (1). However, directly applying RGE or CGE in GDA cannot exhibit satisfactory performance in both single-step and overall complexity (Hu and Fu, 2025). To address this, we combine the framework of random block updates with GDA and smoothed GDA to develop two novel algorithms, called zeroth-order block gradient descent ascent (ZOB-GDA) and zeroth-order block smoothed gradient descent ascent (ZOB-SGDA). Rather than estimating a full gradient at each step, they randomly select a block of coordinates and update the variables using block coordinate-wise gradient estimations (BCGEs). The adoption of the BCGEs effectively controls the bias and variance of gradient estimations to be negligible and thereby accelerates convergence. Moreover, the block size is adjustable to control the number of queries required to construct a single-step gradient.
Best-Known Query Complexity Bounds with Controllable Single-Step Efficiency.
We establish finite-sample guarantees for the proposed algorithms by analyzing the min-max problems (2) in nonconvex-concave settings. The query complexity bounds are summarized in Table 1 and compared with two representative algorithms. Specifically, ZOB-GDA can find an $\epsilon$-stationary point of $f(x,y)$ with a query complexity bound $\mathcal{O}(d/\epsilon^{6})$, which differs from the bound for first-order GDA only by a factor of $d$. Moreover, ZOB-SGDA is shown to have the query complexity bound $\mathcal{O}(d/\epsilon^{4})$, which is consistent with the best available results in the literature for solving deterministic nonconvex-concave problems. Different from Xu et al. (2024), our algorithms benefit from controllable efficiency in single-step gradient estimation, which makes them query-efficient for both single-step and overall complexities. The numerical results demonstrate that our algorithms require significantly fewer queries for both a single step and overall convergence compared to existing methods.
@@ 1.2  Related Work
Here, we review the related work on ZO, zeroth-order GDA, and zeroth-order coordinate/block updates.
@@
ZO has emerged as a prevalent tool to solve black-box problems and found application across machine learning (Liu et al., 2020a; Nguyen and Balasubramanian, 2023), power systems (Hu et al., 2024; Jin et al., 2026), simulation optimization (Xu and Zheng, 2023; Lam and Zhang, 2025), large language models (Malladi et al., 2023; Zhang et al., 2024), etc. ZO originates from the stochastic approximation method in Kiefer and Wolfowitz (1952), where CGE is applied to estimate partial gradients along all dimensions via finite differences of function values. This is inefficient for high-dimensional problems, even if parallel techniques can be applied (Scheinberg, 2022). To improve single-step efficiency, one-point and two-point RGEs have been developed by estimating gradients along randomized directions (Flaxman et al., 2005; Nesterov and Spokoiny, 2017; Lam and Zhang, 2025). Generally, RGE-based algorithms can achieve the same oracle complexities as their first-order counterparts in unconstrained problems, differing by a dimension-dependent factor (Liu et al., 2020a). However, RGEs suffer from large variance in gradient estimation in constrained problems (see Section 2.2 for a detailed discussion). Moreover, most literature considers simple constraints on the input $x$ that can be dealt with by projection operations (Duchi et al., 2015; Yuan et al., 2015; Jin et al., 2026; He et al., 2024), which cannot be used to solve problem (1).
@@
GDA is a classical framework for solving min-max problems and has been extensively studied (Nemirovski, 2004; Nedić and Ozdaglar, 2009; Lin et al., 2020; Xu et al., 2023; Zhang et al., 2020). It is also well-established and widely applied to solve zeroth-order min-max problems of the form (2) (Hu et al., 2024; Zhou et al., 2025). The authors of Liu et al. (2018) applied the two-point RGE to solve a composite optimization problem. Then, the standard zeroth-order GDA was applied to the general min-max problems (Liu et al., 2020b; Wang et al., 2023), while only the query complexity bound of nonconvex-strongly concave cases was established. Several variants of zeroth-order GDA have been developed for convex-concave settings, such as zeroth-order OGDA-RR (Maheshwari et al., 2022) and zeroth-order extra-gradient (Zhou et al., 2025), which can achieve the query complexity bounds of $\mathcal{O}(d^{4}/\epsilon^{2})$ and $\mathcal{O}(d/\epsilon^{2})$ to derive an $\epsilon$-optimal solution, respectively. For nonconvex-concave problems, Nguyen and Balasubramanian (2023) and An et al. (2024) applied RGEs to design algorithms that achieve the query complexity bound of $\mathcal{O}(d/\epsilon^{6})$ to derive an $\epsilon$-stationary/KKT point. In addition, Xu et al. (2024) proposed combining alternating gradient projection with CGEs to solve a min-max problem with a complexity bound of $\mathcal{O}(d/\epsilon^{4})$ (matching the best available bound) to obtain an $\epsilon$-stationary point.
@@
The framework of coordinate/block updates is widely adopted in first-order optimization (Nesterov and Stich, 2017; Latafat et al., 2019). The core idea is to apply the partial gradients along a subset of full dimensions to update the iterates. The applications of coordinate/block updates in ZO mainly lie in unconstrained problems (Lian et al., 2016; Cai et al., 2021), where only the coordinate/block gradients along a subset are estimated using coordinate/block CGEs or RGEs at each step. Their extension to constrained problems, however, remains relatively underexplored. In Shanbhag and Yousefian (2021), the RGE was combined with zeroth-order block updates and projected gradient descent to solve a stochastic constrained problem. Moreover, in He et al. (2024) and Jin et al. (2026), a cyclic zeroth-order block coordinate descent method and a randomized zeroth-order coordinate descent method were proposed, respectively, to solve the deterministic constrained problems and achieve complexity bounds proportional to $\epsilon^{-2}$ for nonconvex optimization. However, all these methods require the constraint set to be projection-friendly and coordinate/block-structured, which is usually too restrictive in practical problems and inapplicable to non-analytical constraint sets (such as in our problem (1)).
@@ 2  Preliminaries
For a positive integer $n$, we denote $n]:=\{1,2,\cdots,n\}$. For a vector $x\in\mathbb{R}^{d_{x}}$, denote $(x)_{i}$ as its $i$th entry. For a differentiable function $h(x):\mathbb{R}^{d_{x}}\to\mathbb{R}$, denote $\nabla h(x)$ as its gradient at $x$ and $\nabla_{i}h(x),i\in[d_{x}]$ as the partial gradient along the $i$th dimension. Similarly, for a differentiable function $f(x,y):\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$, denote the partial gradient w.r.t. $x$ (and $y$) by $\nabla_{x}f(x,y)$ (and $\nabla_{y}f(x,y)$). Without further specification, $\|\cdot\|$ denotes the $\ell_{2}$-norm in Euclidean space. The Euclidean projection operator onto a closed convex set $\mathcal{X}$ is denoted by $\mathcal{P}_{\mathcal{X}}[\cdot]$.